# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v2

@@ VI-B8 Topological Nanowire Qubits (Majorana Qubits)
VI-C Quantum Circuit Mapping
@@
Combining one-qubit rotation gates and two-qubit Controlled-NOT (CNOT) gates can realize any quantum circuit. However, CNOT gates cannot be applied to all pairs of qubits in the currently available NISQ quantum devices. Therefore, the circuit must be transformed into an equivalent circuit that does not violate the limitation of the device. The transformation often requires additional gates that limit the usability of the device [158]. The quality of the circuit mapping can be assessed via different metrics, such as the resulting circuit’s (two-qubit) gate count, depth, or the expected fidelity [159].
@@
VI-D Quantum Simulators and Primitives
@@
Several software simulators have been released by major companies involved in QML research. While Table IV summarizes the key features of the simulators, the following list provides key details of these tools.
  • IBM Qiskit [47] is an open-source software development kit for quantum systems at the level of circuits, pulses, and algorithms. With the help of domain-specific APIs, Qiskit builds a software stack that makes it easy for users to use quantum computers, as it allows practitioners to easily design experiments and applications and run them on classical simulators or real quantum computers. An overview of the main characteristics of Qiskit is illustrated in Figure 32. It comprises the following components:
@@
  • Xanadu PennyLane [48] is the leading tool for programming quantum computers. It is a cross-platform Python library that enables quantum differentiable programming and integration with ML tools to train a quantum computer like you would train a DNN. PennyLane also supports a comprehensive set of features, hardware, simulators, and community-led resources that enable users of all levels to build, optimize, and deploy quantum-classical applications easily. An overview of its main features is depicted in Figure 33.
@@
  • Xanadu Strawberry Fields [160] is a full-stack Python library for designing quantum algorithms and executing them directly on photonic quantum computers such as Xanadu’s next-generation quantum hardware.
  • Rigetti PyQuil [161, 162] is a Python library for writing and running quantum programs using Quantum instruction language (Quil) programs. PyQuil allows developers to easily generate Quil programs from quantum gates and classical operations, compile and simulate these programs using the Quil compiler (Quilc) and the Quantum Virtual Machine (QVM), or execute them on a Rigetti quantum processor. Quilc is an optimizing compiler for gate-based quantum programs, capable of performing circuit transformations to produce optimal circuit implementations for a specific Rigetti processor. Such optimizations allow developers to write programs faster while preserving (or improving) their execution fidelity on a given hardware system. Quil is built and executed using the Forest SDK, a software tool that allows users to write quantum programs in Quil.
  • Quantinuum (Honeywell) TKET [163] is an advanced software development kit, accessible through the PyTKET Python package, for the creation and execution of programs for gate-based quantum computers. It is platform-inclusive, and its state-of-the-art circuit optimization routines allow users to extract as much power as possible from any of today’s NISQ devices.
  • Google Cirq [164] is an open-source framework for programming quantum computers, whose software library can be used for writing, manipulating, optimizing, and running quantum circuits on quantum computers and quantum simulators. Cirq provides valuable abstractions for dealing with today’s NISQ computers, where hardware details are vital to achieving state-of-the-art results.
  • Google OpenFermion [165] is an open-source library for analyzing and compiling quantum algorithms to simulate fermionic systems, including quantum chemistry. The package provides efficient data structures for representing fermionic operators and fermionic circuit primitives for their execution on quantum devices. Plugins to OpenFermion provide users with efficient and low-overhead means of translating electronic structure calculations into quantum circuit calculations.
  • Google TensorFlow Quantum (TFQ) [166] is a QML library for rapid prototyping hybrid quantum-classical ML models. Research in quantum algorithms and applications can leverage Google’s TFQ, all from within TensorFlow. It integrates quantum computing algorithms and logic designed in Cirq and provides quantum computing primitives compatible with existing TensorFlow APIs, along with high-performance quantum circuit simulators.
  • D-Wave Leap [167] is a programming model and execution framework to execute quantum and hybrid workloads effectively.
VI-E Supporting Infrastructure Tools and Hardware
@@
  • qBraid [168] is a cloud-based quantum computing platform designed to accelerate the pace of progress in the field. It offers access to the ecosystem of quantum software, hardware, and an entire suite of application-specific quantum algorithms, helping software engineers run programs on quantum computers available from AWS, IBM, Rigetti, IonQ, and QuEra within minutes. qBraid Lab is a web-based IDE interface (through JupiterLab) that provides software tools for developers and access to quantum hardware. It also includes CLI, a Command Line Interface for interacting with all parts of the qBraid platform.
@@
  • IonQ Quantum Cloud [169] provides access compatible with most Quantum SDKs with their continuous effort to extend the support to newer tools. The cloud provides access to all IonQ’s quantum processors and supports noisy and ideal simulators.
@@
  • Amazon Braket by AWS [170] is a fully managed quantum computing service designed to speed up scientific research and software development for quantum computing. Its support is useful for experimenting with quantum computing algorithms, testing different quantum hardware, and building and developing quantum software. It provides access to different quantum computers and simulators using a consistent collection of development tools. Users can build quantum projects on a trusted cloud with affordable pricing and management controls for both quantum and classical workloads. It allows faster execution of hybrid quantum-classical algorithms with priority access to quantum computers and no classical infrastructure to manage.
@@
  • Microsoft Azure Quantum [171] provides quantum cloud access equipped with open-source development toolkits and environments for most leading quantum technologies with runtime metrics. Microsoft is adopting topological qubits as their qubit modality for developing their quantum computers. Their aim behind using these qubits directly corresponds with the long-term scalability of quantum computers. Since the topological qubits can survive at regular room temperatures, their usage improves the systems’ stability compared to other qubit technologies. The resource estimator tool provided by Azure monitors the quantum system’s runtime statistics. It estimates the resources allocated by the quantum code and refines the solution to a reasonable extent. This tool calculates the number of logical and physical qubits and runtime requirements to execute quantum applications on future-scaled quantum computers. Additional features, like determining the number of physical qubits required by circuits and evaluating the differences across qubit technologies, facilitate the design of quantum solutions with varying technologies and refining them to scaled versions for FTQC.
@@
  • Zapata Orquestra [172] is a platform to develop and deploy generative AI applications. It allows rapid prototyping, benchmarking, and executing workflows across quantum and classical computing resources. Their objectives focus on prioritizing use cases that can deliver a near-term financial impact and build customer first-mover capabilities, ultimately leading to solving computationally complex problems.
@@
  • Xanadu Cloud [173] is an interface that allows the simulation and execution of programs on quantum photonic hardware. Users can develop quantum applications and program quantum computers with PennyLane’s library of tools, demonstrations, tutorials, and community support forums.
VII Applications
@@
VII-A Molecular Simulation: Drug Design
@@
QC’s primary value for pharmaceuticals lies in the R&D phase of drug design [174]. QML can help drug developers build predictive models to identify which molecules will likely be effective drug candidates. These models can filter out less promising candidates early in the drug development process, saving resources and time [175]. A practical use-case of this application is represented by material modeling [176], which can be efficiently implemented with QML.
VII-B Quantum Sensing
Quantum Sensing is an advanced sensor technology that detects changes in motion and electromagnetic fields by collecting data at the atomic level [177].
@@ VII-B Quantum Sensing
As witnessed by the increasing number of startups in this field, quantum sensors are finding their way from laboratories to the real world. The atomic length scale of quantum sensors and their coherence properties enable unprecedented spatial resolution and sensitivity. Biomedical applications ranging from brain imaging to single-cell spectroscopy [178] could benefit from these quantum technologies, but evaluating the potential impact of the techniques is not trivial.
@@
Metrology is another applicable domain for quantum sensing [179]. The current interests in this direction demand compact and susceptible measurement systems. Integrating electronics with quantum sensing technology should be one of the viable areas of interest for NISQ devices.
VII-C Quantum Cryptography
@@
This application domain entirely focuses on creating new and improved cryptography schemes that ensure security against attacks and threats by quantum breach technology. Proactively taking measures against the utilization of quantum technology for breaking such schemes is referred to as Quantum Safe cryptography. Keeping that in mind, currently used encryption methods for many critical communication and transaction processes use the Rivest–Shamir–Adleman (RSA) scheme [180]. It is fundamentally based on the keys generated from prime factors of very large numbers. Although any classical computer can find prime numbers of a given small number given the understanding of complexity theory, it has been proven that for large numbers, the complexity of finding its prime numbers becomes difficult to intractable, i.e., solvable only in indefinite time. This particular property of the problem with classical computers made it an obvious choice to use the RSA as an encryption method. Something that would require thousands of years to solve with the available technology would raise no harm. However, the works of Peter Shor [181, 182] showed the capability of quantum computers to obtain solutions to problems that were presumably be indefinitely solved in the classical domain. This single instance of challenges to established security claims opens new avenues, driven by the need to develop improved and robust cryptography methods that make the communication Quantum Safe [183, 184]. Instead of completely changing a classical algorithm into quantum, the focus is on adding quantum modules to existing classical cryptography techniques. The advantage of this strategy is that while the existing classical techniques remain intact, applying the quantum methods improves their efficiency and speed [185]. This proved to be a very useful domain of potential applications for QC that builds over the existing schemes and makes them robust and efficient rather than re-inventing the whole process.
VII-D Financial Modeling
@@
By enabling more accurate and complex financial data modeling, improving risk management, and optimizing investment portfolios, QML could lead to better and informed decision-making with higher returns for investors. It can also enable financial services organizations to re-engineer operational processes [186]. QML’s specific use cases for financial services can be classified into three main categories: targeted predictions, trading optimizations, and risk profiling. Quantum technology computes more efficiently than classical computing when generating probability distributions, mapping data, testing samples, and iterating. These properties align well with the financial modeling since they can help enhance investment gains, reduce capital requirements, open new investment opportunities, and improve the identification and management of risk and compliance. The simplest instance of practical implementation of QML for finance can include a task of portfolio optimization using numerous quantum optimization algorithms, fundamentally derived from the parameterized quantum circuits. If not at the current stage, ultimately such tasks will benefit in terms of compute speed and processing information capacity for solutions using QML techniques.
VII-E Logistics Optimization
@@
VII-F Environmental, Social, and Governance (ESG)
@@
VIII Conclusion and Road Ahead
@@
  • Scalability: Achieving scalability in quantum computing involves moving beyond single quantum chips to develop multi-chip devices with efficient interconnection and communication systems. Companies like IBM and Honeywell are making strides toward large-scale quantum devices, as shown in Figure 14, but significant challenges remain. The primary focus should be on minimizing error rates and enhancing system reliability. As suggested in [32], reducing error rates is a top priority for making quantum computation technology both impactful and widely adopted. Innovations in error correction and fault-tolerant quantum computing will be crucial for scaling up quantum systems.
@@ Appendix A Complexity Theory
This section provides an overview of complexity classes. A promise problem is a decision problem where the input is to be selected from all possible input strings. A problem with an explicitly defined structure or characteristics of input is defined in advance, making the problem environment-entrusted with certain types of promised input with no instance of input sent in unexpectedly or out of definition. All promise problems are assigned to a complexity class suitable to their nature and requirements. Formally, complexity classes are a group of computational problems that have similar resource-based complexity [187]. In classical computing, there exist the following classes:
@@ Appendix B Quantum Technical Concepts and Definitions
B-A Hilbert Space
In Quantum mechanics the state of a physical system is represented by a vector in the Hilbert space, which is a complex vector space with an inner product. In the finite-dimensional complex vector spaces that come up in quantum computation and quantum information, the Hilbert space is exactly the same thing as an inner product space (both terms can be used interchangeably) [^2] For infinite dimensions, the Hilbert space satisfies additional technical restrictions above and beyond inner product spaces [36]..
B-B State Space and State Vector
The Hilbert space is associated with any isolated physical system, known as the state space of the system [36]. The system is completely described by its state vector, which is a unit vector in the system’s state space.
B-C Schrodinger’s Equation
The Schrödinger’s equation (see Equation 4) describes the evolution of the state $\psi$ of a closed quantum system over time $t$ [36].

### v3

@@ 6.2.8 Topological Nanowire Qubits (Majorana Qubits)
6.3 Quantum Circuit Mapping
@@
Combining one-qubit rotation gates and two-qubit Controlled-NOT (CNOT) gates can realize any quantum circuit. However, CNOT gates cannot be applied to all pairs of qubits in the currently available NISQ quantum devices. Therefore, the circuit must be transformed into an equivalent circuit that does not violate the limitation of the device. The transformation often requires additional gates that limit the usability of the device (Itoko et al., 2020). The quality of the circuit mapping can be assessed via different metrics, such as the resulting circuit’s (two-qubit) gate count, depth, or the expected fidelity (Wille & Burgholzer, 2023).
@@
6.4 Quantum Simulators and Primitives
@@
Several software simulators have been released by major companies involved in QML research. While Table 4 summarizes the key features of the simulators, the following list provides key details of these tools.
  • IBM Qiskit (Qiskit contributors, 2023) is an open-source software development kit for quantum systems at the level of circuits, pulses, and algorithms. With the help of domain-specific APIs, Qiskit builds a software stack that makes it easy for users to use quantum computers, as it allows practitioners to easily design experiments and applications and run them on classical simulators or real quantum computers. An overview of the main characteristics of Qiskit is illustrated in Figure 33. It comprises the following components:
@@
  • Xanadu PennyLane (Bergholm et al., 2018) is the leading tool for programming quantum computers. It is a cross-platform Python library that enables quantum differentiable programming and integration with ML tools to train a quantum computer like you would train a DNN. PennyLane also supports a comprehensive set of features, hardware, simulators, and community-led resources that enable users of all levels to build, optimize, and deploy quantum-classical applications easily. An overview of its main features is depicted in Figure 34.
@@
  • Xanadu Strawberry Fields (Killoran et al., 2019) is a full-stack Python library for designing quantum algorithms and executing them directly on photonic quantum computers such as Xanadu’s next-generation quantum hardware.
  • Rigetti PyQuil (Smith et al., 2016; Karalekas et al., 2020) is a Python library for writing and running quantum programs using Quantum instruction language (Quil) programs. PyQuil allows developers to easily generate Quil programs from quantum gates and classical operations, compile and simulate these programs using the Quil compiler (Quilc) and the Quantum Virtual Machine (QVM), or execute them on a Rigetti quantum processor. Quilc is an optimizing compiler for gate-based quantum programs, capable of performing circuit transformations to produce optimal circuit implementations for a specific Rigetti processor. Such optimizations allow developers to write programs faster while preserving (or improving) their execution fidelity on a given hardware system. Quil is built and executed using the Forest SDK, a software tool that allows users to write quantum programs in Quil.
  • Quantinuum (Honeywell) TKET (Qua, ) is an advanced software development kit, accessible through the PyTKET Python package, for the creation and execution of programs for gate-based quantum computers. It is platform-inclusive, and its state-of-the-art circuit optimization routines allow users to extract as much power as possible from any of today’s NISQ devices.
  • Google Cirq (Developers, 2023) is an open-source framework for programming quantum computers, whose software library can be used for writing, manipulating, optimizing, and running quantum circuits on quantum computers and quantum simulators. Cirq provides valuable abstractions for dealing with today’s NISQ computers, where hardware details are vital to achieving state-of-the-art results.
  • Google OpenFermion (McClean et al., 2020) is an open-source library for analyzing and compiling quantum algorithms to simulate fermionic systems, including quantum chemistry. The package provides efficient data structures for representing fermionic operators and fermionic circuit primitives for their execution on quantum devices. Plugins to OpenFermion provide users with efficient and low-overhead means of translating electronic structure calculations into quantum circuit calculations.
  • Google TensorFlow Quantum (TFQ) (Broughton et al., 2020) is a QML library for rapid prototyping hybrid quantum-classical ML models. Research in quantum algorithms and applications can leverage Google’s TFQ, all from within TensorFlow. It integrates quantum computing algorithms and logic designed in Cirq and provides quantum computing primitives compatible with existing TensorFlow APIs, along with high-performance quantum circuit simulators.
  • D-Wave Leap (DWa, ) is a programming model and execution framework to execute quantum and hybrid workloads effectively.
6.5 Supporting Infrastructure Tools and Hardware
@@
  • qBraid (qBr, ) is a cloud-based quantum computing platform designed to accelerate the pace of progress in the field. It offers access to the ecosystem of quantum software, hardware, and an entire suite of application-specific quantum algorithms, helping software engineers run programs on quantum computers available from AWS, IBM, Rigetti, IonQ, and QuEra within minutes. qBraid Lab is a web-based IDE interface (through JupiterLab) that provides software tools for developers and access to quantum hardware. It also includes CLI, a Command Line Interface for interacting with all parts of the qBraid platform.
@@
  • IonQ Quantum Cloud (Ion, ) provides access compatible with most Quantum SDKs with their continuous effort to extend the support to newer tools. The cloud provides access to all IonQ’s quantum processors and supports noisy and ideal simulators.
@@
  • Amazon Braket by AWS (Amazon Web Services, 2020) is a fully managed quantum computing service designed to speed up scientific research and software development for quantum computing. Its support is useful for experimenting with quantum computing algorithms, testing different quantum hardware, and building and developing quantum software. It provides access to different quantum computers and simulators using a consistent collection of development tools. Users can build quantum projects on a trusted cloud with affordable pricing and management controls for both quantum and classical workloads. It allows faster execution of hybrid quantum-classical algorithms with priority access to quantum computers and no classical infrastructure to manage.
@@
  • Microsoft Azure Quantum (Mic, ) provides quantum cloud access equipped with open-source development toolkits and environments for most leading quantum technologies with runtime metrics. Microsoft is adopting topological qubits as their qubit modality for developing their quantum computers. Their aim behind using these qubits directly corresponds with the long-term scalability of quantum computers. Since the topological qubits can survive at regular room temperatures, their usage improves the systems’ stability compared to other qubit technologies. The resource estimator tool provided by Azure monitors the quantum system’s runtime statistics. It estimates the resources allocated by the quantum code and refines the solution to a reasonable extent. This tool calculates the number of logical and physical qubits and runtime requirements to execute quantum applications on future-scaled quantum computers. Additional features, like determining the number of physical qubits required by circuits and evaluating the differences across qubit technologies, facilitate the design of quantum solutions with varying technologies and refining them to scaled versions for FTQC.
@@
  • Zapata Orquestra (Zap, ) is a platform to develop and deploy generative AI applications. It allows rapid prototyping, benchmarking, and executing workflows across quantum and classical computing resources. Their objectives focus on prioritizing use cases that can deliver a near-term financial impact and build customer first-mover capabilities, ultimately leading to solving computationally complex problems.
@@
  • Xanadu Cloud (Xan, ) is an interface that allows the simulation and execution of programs on quantum photonic hardware. Users can develop quantum applications and program quantum computers with PennyLane’s library of tools, demonstrations, tutorials, and community support forums.
7 Applications
@@
7.1 Molecular Simulation: Drug Design
@@
QC’s primary value for pharmaceuticals lies in the R&D phase of drug design (Epifanovsky et al., 2021). QML can help drug developers build predictive models to identify which molecules will likely be effective drug candidates. These models can filter out less promising candidates early in the drug development process, saving resources and time (Santagati et al., 2023). A practical use-case of this application is represented by material modeling (Lourenço et al., 2023), which can be efficiently implemented with QML.
7.2 Quantum Sensing
Quantum Sensing is an advanced sensor technology that detects changes in motion and electromagnetic fields by collecting data at the atomic level (Degen et al., 2017).
@@ 7.2 Quantum Sensing
As witnessed by the increasing number of startups in this field, quantum sensors are finding their way from laboratories to the real world. The atomic length scale of quantum sensors and their coherence properties enable unprecedented spatial resolution and sensitivity. Biomedical applications ranging from brain imaging to single-cell spectroscopy (Aslam et al., 2023) could benefit from these quantum technologies, but evaluating the potential impact of the techniques is not trivial.
@@
Metrology is another applicable domain for quantum sensing (Ohshima, 2022). The current interests in this direction demand compact and susceptible measurement systems. Integrating electronics with quantum sensing technology should be one of the viable areas of interest for NISQ devices.
7.3 Quantum Cryptography
@@
This application domain entirely focuses on creating new and improved cryptography schemes that ensure security against attacks and threats by quantum breach technology. Proactively taking measures against the utilization of quantum technology for breaking such schemes is referred to as Quantum Safe cryptography. Keeping that in mind, currently used encryption methods for many critical communication and transaction processes use the Rivest–Shamir–Adleman (RSA) scheme (Rivest et al., 1978). It is fundamentally based on the keys generated from prime factors of very large numbers. Although any classical computer can find prime numbers of a given small number given the understanding of complexity theory, it has been proven that for large numbers, the complexity of finding its prime numbers becomes difficult to intractable, i.e., solvable only in indefinite time. This particular property of the problem with classical computers made it an obvious choice to use the RSA as an encryption method. Something that would require thousands of years to solve with the available technology would raise no harm. However, the works of Peter Shor (Shor, 1994; 1997) showed the capability of quantum computers to obtain solutions to problems that were presumably be indefinitely solved in the classical domain. This single instance of challenges to established security claims opens new avenues, driven by the need to develop improved and robust cryptography methods that make the communication Quantum Safe (Campagna et al., 2015; Stebila & Mosca, 2016). Instead of completely changing a classical algorithm into quantum, the focus is on adding quantum modules to existing classical cryptography techniques. The advantage of this strategy is that while the existing classical techniques remain intact, applying the quantum methods improves their efficiency and speed (Bozzio et al., 2022). This proved to be a very useful domain of potential applications for QC that builds over the existing schemes and makes them robust and efficient rather than re-inventing the whole process.
7.4 Financial Modeling
@@
By enabling more accurate and complex financial data modeling, improving risk management, and optimizing investment portfolios, QML could lead to better and informed decision-making with higher returns for investors. It can also enable financial services organizations to re-engineer operational processes (Orús et al., 2019). QML’s specific use cases for financial services can be classified into three main categories: targeted predictions, trading optimizations, and risk profiling. Quantum technology computes more efficiently than classical computing when generating probability distributions, mapping data, testing samples, and iterating. These properties align well with the financial modeling since they can help enhance investment gains, reduce capital requirements, open new investment opportunities, and improve the identification and management of risk and compliance. The simplest instance of practical implementation of QML for finance can include a task of portfolio optimization using numerous quantum optimization algorithms, fundamentally derived from the parameterized quantum circuits. If not at the current stage, ultimately such tasks will benefit in terms of compute speed and processing information capacity for solutions using QML techniques.
7.5 Logistics Optimization
@@
7.6 Environmental, Social, and Governance (ESG)
@@
8 Conclusion and Road Ahead
@@
  • Scalability: Achieving scalability in quantum computing involves moving beyond single quantum chips to develop multi-chip devices with efficient interconnection and communication systems. Companies like IBM and Honeywell are making strides toward large-scale quantum devices, as shown in Figure 14, but significant challenges remain. The primary focus should be on minimizing error rates and enhancing system reliability. As suggested in (Preskill, 2018), reducing error rates is a top priority for making quantum computation technology both impactful and widely adopted. Innovations in error correction and fault-tolerant quantum computing will be crucial for scaling up quantum systems.
@@ Appendix A Complexity Theory
This section provides an overview of complexity classes. A promise problem is a decision problem where the input is to be selected from all possible input strings. A problem with an explicitly defined structure or characteristics of input is defined in advance, making the problem environment-entrusted with certain types of promised input with no instance of input sent in unexpectedly or out of definition. All promise problems are assigned to a complexity class suitable to their nature and requirements. Formally, complexity classes are a group of computational problems that have similar resource-based complexity (Babai et al., 1986). In classical computing, there exist the following classes:
@@ Appendix B Quantum Technical Concepts and Definitions
B.1 Hilbert Space
In Quantum mechanics the state of a physical system is represented by a vector in the Hilbert space, which is a complex vector space with an inner product. In the finite-dimensional complex vector spaces that come up in quantum computation and quantum information, the Hilbert space is exactly the same thing as an inner product space (both terms can be used interchangeably) [^2] For infinite dimensions, the Hilbert space satisfies additional technical restrictions above and beyond inner product spaces (Nielsen & Chuang, 2010b)..
B.2 State Space and State Vector
The Hilbert space is associated with any isolated physical system, known as the state space of the system (Nielsen & Chuang, 2010b). The system is completely described by its state vector, which is a unit vector in the system’s state space.
B.3 Schrodinger’s Equation
The Schrödinger’s equation (see Equation 4) describes the evolution of the state $\psi$ of a closed quantum system over time $t$ (Nielsen & Chuang, 2010b).